OpenStudy (anonymous):

simplify w^8x^4w^10x^5

OpenStudy (anonymous):

When multiplying like terms with different exponents you add the exponents. Since you have two different terms here, x and w, you will multiply the w terms together and the x terms together. So for w^8 * w^10 = w^(8+10) = w^18. Using the same method, you should be able to simplify the x terms.

OpenStudy (anonymous):

wait i dont get it so what is the answer?

OpenStudy (anonymous):

I assumed you wanted to know how to do the problem, not just be given an answer. It would be far more helpful if you learned how the properties of exponents work.

OpenStudy (anonymous):

true but im confused i learn math slow? slow learner

OpenStudy (anonymous):

I understand. So say you had the following expression: t^3*t^4. When multiplying like terms, which in this case is t, with different exponents, all you need to do is add the exponents. so t^3*t^4 =t^(3+4) = t^7. So going back to your original equation w^8*x^4w^10*x^5. First you must group like terms. So your original equation becomes: w^8*w^10* x^4*x^5. You would then combine like terms by adding exponents, so you'd get w^(8+10)*x^(4+5).